# Soft Masked Transformer for Point Cloud Processing with Skip Attention-Based Upsampling

### v1

@@ Methods
Figure 6: Visualization of semantic segmentation results on SemanticKITTI (each scene is 50m by 50m, centered around the LiDAR). To accentuate the disparity between predictions and ground truth, we colour right/wrong predictions in gray/red colour, respectively. Notice that there are very few red points i.e. very few wrong predictions.
Results: Table IV presents the outcomes of our network alongside results from well-established methods on the SemanticKITTI dataset. Our approach demonstrates commendable performance, achieving 74.9% mIoU. Notably, compared to the cutting-edge projection-based method RangFormer, which is pre-trained on other datasets, our method exhibits superior performance (+1.6%) without any pre-training. Furthermore, our proposed method surpasses voxel partitioning and 3D convolution-based techniques, such as Cylinder3D, 2DPASS, and PVKD. Importantly, our model showcases a remarkable understanding of certain small object categories, including poles, traffic signs and motorcyclists. Fig. 6 shows visualizations of our results on SemanticKITTI validation set.
@@
We present the results for 20 classes of interest in the SWAN test frames in Table V. We compare our results to PointNet++ [22], PointConv [13], Cylinder3D [68], PolarNet [67], $\psi$-CNN [75] and SAT3D [72]. As depicted in Table V, the mIoU values for these compared methods on the SWAN dataset are lower than those on the semanticKITTI dataset. This discrepancy can be attributed to the heightened complexity of the scenes in the SWAN dataset which was captured in dense central business district are of the city of Perth, Australia. Notably, on this dataset, our method demonstrates the best performance with a remarkable improvement of +4.2% in mIoU compared to the nearest competitor SAT3D. Our method is generally able to show remarkable prediction accuracy towards some small objects, including light poles, traffic sign and pedestrians.
IV-C Object Classification
Datasets: We evaluate SMTransformer on the synthetic data ModelNet40[76], and real-world data ScanobjectNN[77]. The ModelNet40 comprises 12,311 CAD models from 40 categories and is divided into 9,843 training and 2,468 test models. Each sample has about 10,000 points and the features contain coordinates and normals. The ScanObjectNN dataset comprises 15,000 objects categorized into 15 classes, selected from ScanNetv2. In contrast to the synthetic ModelNet40 objects, these objects exhibit occlusion, background noise, deformed geometric shapes, and non-uniform surface density, presenting a more challenging scenario. Our experiments are conducted on its most challenging perturbed variant, denoted as PB__\__T50__\__RS. Here, we uniformly sample 1024 points from each model and only use their $(X,Y,Z)$ coordinates as input. We follow the data augmentation used in PointNext[47].
Network Configurations: We employ identical network configurations for both ModelNet40 and ScanObjectNN. Throughout the training, we utilize the SGD optimizer with a momentum of 0.9 and an initial learning rate set to 0.1, training the model for 350 epochs using a batch size of 32. To dynamically adapt the learning rate, we implement cosine annealing, adjusting it when it decreases to 0.001 and applying a dropout ratio of 0.4.
TABLE VI: Classification results on ModelNet40 and ScanObjectNN dataset. ‘xyz’ and ‘n’ represent coordinates and normal vector. ‘K’ stands for one thousand and ‘PN++’ for PointNet++. Our network achieves the best overall accuracy.
Methods Input #Points OA(%)
ModelNet40 ScanObjectNN
PointWeb[30] xyz, n 1K 92.3 -
PointConv[13] xyz, n 1K 92.5 -
SpiderCNN[12] xyz, n 5K 92.4 -
KPConv[10] xyz 7K 92.9 -
PointASNL[21] xyz, n 1K 93.2 -
PRANet[78] xyz 2K 93.7 82.1
RS-CNN[79] xyz 1K 93.6 -
PointNet[24] xyz 1K 89.2 68.2
PointNet++[25] xyz 1K 90.7 77.9
DGCNN[33] xyz 1K 92.9 78.2
PointCNN[36] xyz 1K 92.2 78.5
BGA-DGCN[77] xyz 1K - 79.9
BGA-PN++[77] xyz 1K - 80.2
PointASNL[21] xyz 1K 92.9 -
PRANet[78] xyz 1K 93.2 81.0
PointTransformer[23] xyz 1K 93.7
PointMLP[28] xyz 1K 94.1 85.4
PointTransformerV2[22] xyz 1K 94.2
DANet[80] xyz 1K 93.6
LCPformer[53] xyz 1K 93.6
PointNext[47] xyz 1K 93.2 87.7
PointVector[47] xyz 1K 87.8
SMTransformer(ours) xyz 1K 94.2 88.0
Results: We compare our method with representative state-of-the-art methods in Table VI using the overall accuracy (OA) metric. For better comparison, we also show the input data type and the number of input points for each method. Our network achieves the best performance of 94.2% OA on ModelNet40 and the best performance of 88.0% OA on ScanObjectNN.
On ModelNet40, our network surpasses the classical local point convolution KPConv by 1.3%, even though KPConv uses 7,000 input points while our network uses only 1,024 points. In comparison to the previous state-of-the-art MLP-based method (e.g., PointNext), our network outperforms it by 1%. Furthermore, compared to the classical vector attention-based transformers, our network outperforms the Point Transformer by 0.6% and achieves similar competitive results 94.2% as the Point Transformer V2 in terms of OA. Compared to the scalar attention-based methods, our network outperforms the PointANSL and LCPFormer by 1% and 0.6%, respectively.
On ScanObjectNN, our network achieves the state-of-the-art overall accuracy (OA) of 88.0%. Specifically, when compared to MLP-based methods, our network outperforms PointMLP, PointNext, and PointVector by 2.6%, 0.3%, and 0.2% in terms of OA. This superior performance on a real-world dataset highlights the suitability of our method for practical applications.
IV-D Ablation Studies
@@
We conduct ablation studies on the S3DIS dataset to demonstrate the effectiveness of SMTransformer, SAUB and Shared Position Encoding.
1) Effect of Various Components: Table VII displays the influence of our introduced modules (i.e. SMTransformer and SAUB) and strategy (i.e. shared position encoding). Case I is the baseline Point Transformer which does not include any of our modules. Cases II to IV systematically incorporate each of our proposed components, progressively enhancing the baseline result to reach 73.4%. Case V employs the shared position encoding strategy while maintaining performance similar to the unshared position encoding strategy. This indicates that point clouds with the same resolution could share the same position encoding information without sacrificing accuracy.
TABLE VII: Effect of various components on semantic segmentation (S3DIS Area-5).‘ ${}^{\dagger}$’ donates the Soft Masked Transformer with position encoding used in the Point Transformer[23]. ‘EPE’ denotes the proposed enhanced position encoding. ‘SAUB’ donates the proposed skip attention-based up-sampling block. ‘SPE’ is the shared position encoding strategy. ‘Para.’ donates the network parameters.
Case ${}^{\dagger}$ EPE SAUB SPE mIoU(%) mAcc(%) Para.(M)
I 70.6 76.5 7.8
II 72.0 77.8 8.4
III 72.9 78.2 9.0
IV 73.3 78.8 10.3
V 73.4 78.9 7.8
2) SPE versus Unshared Position Encoding (USPE): We conducted a comparative analysis between a network employing shared position encoding and another utilizing unshared position encoding while maintaining consistency in the remaining network configuration. The findings, as presented in Table VIII, highlight the noteworthy advantages of the shared position encoding strategy. The network with shared position encoding not only delivers a superior performance of 73.4% mIoU, but also boasts efficiency gains with fewer parameters (7.8 million) and shorter training time (16 hours).
TABLE VIII: Comparative Analysis of Network Performance with Shared and Unshared Position Encoding.
Case mIoU(%$\uparrow$ Para.(M)$\downarrow$ Training time(h)$\downarrow$
SMTransformer + USPE 73.3 10.3 24
SMTransformer + SPE 73.4 7.8 16
@@
3) Soft Mask versus Hard Mask: To demonstrate the effectiveness and versatility of the soft mask, we conduct a comparison with the hard mask. The hard mask, represented by binary mask, can be expressed as,
@@
4) SAUB versus Classic Up-sampling: To prove the effectiveness of our proposed skip attention-based up-sampling block (SAUB), we compare it with two types of up-sampling blocks including the transition up-sampling block (TUB)[23] and grid unpooling block (GUB)[22]. TUB consists of one projection layer, one interpolation, and one addition operation. GUB consists of one projection layer, one grid unpooling operation and one addition operation. The addition operation connects the features from the encoding and decoding layers. Table IX presents the performance of our network with various up-sampling blocks. The network employing the SAUB achieves the best performance with mIoU, mAcc, and OA values of 73.4%, 78.9%, and 91.8%, respectively, surpassing the performance of the network using the TUB by a significant margin.
TABLE IX: Segmentation performance of our model on S3DIS area 5 with different up-sampling blocks. TUB: transition up-sampling block, GUB: grid unpooling block, SAUB: skip-attention-based up-sampling block.
@@
SAUB(ours) 73.4 78.9 91.8
IV-E Robustness Analysis
1) Robustness to Density: We compare the robustness of our model to inter- and intra-point cloud density with several typical baselines such as PointNet [24], PointNet++ [25], DGCNN[33], classical convolutional network such as PointConv [13], RS-CNN[79], DANet[80] and attention network PointASNL[21]. For a fair comparison, all the networks are trained on modelnet40_normal_resampled dataset[76] with 1024 points using only coordinates as the input. To showcase the robustness across inter-point clouds with varying densities, we utilize downsampled points of 512, 256, 128, and 64 as input to the trained model. To evaluate the robustness of intra-point cloud with various densities, we divide the 1024 points into four equal parts along the $X$ coordinate according to the point number, and then we randomly sample 128 points from each part in sequence. This generates the test samples with 896, 768, 640 and 512 points, respectively. The results are shown in Fig. 8. Our SMTransformer shows significantly superior robustness, surpassing existing approaches for both inter and intra-point cloud variations.
@@ IV-E Robustness Analysis
2) Robustness to Transformation: To demonstrate the robustness of our SMTransformer, we evaluate its performance on S3DIS and ModelNet40 under a variety of perturbations in the test data, including permutation, translation, scaling and jitter. As shown in Table X, on S3DIS, Point Transformer and Point Vector have a huge performance drop on scaling transformation. Our method exhibits remarkable stability across diverse transformations. Particularly noteworthy is its stable performance even amidst a 0.2 translation along the $X$, $Y$, and $Z$ axes and jitter. All methods are invariant to permutations. In terms of sensitivity to point scaling, SMTransformer performs relatively better when the scaling range is decreased. Our method achieves the best accuracy under all transformations on both segmentation and classification datasets.
TABLE X: Robustness study for random point permutations, translation of $\pm$ 0.2 in $X,Y,Z$ axis, scaling ($\times$0.8,$\times$1.2) and jittering. Note that this ablation study is without test time augmentation.
@@
3) Robustness to Noise: To assess the robustness of SMTransformer to noise, we conducted experiments using the PB_T50_RS variant of ScanObjectNN dataset, measuring the performance with and without background noise (denoted as ‘obj_bg’ and ‘obj_nobg’ respectively). Table XI presents a comparative analysis between our model and several baselines from [77]. We observe that the overall accuracy of all networks diminishes when trained and tested under conditions involving background noise. However, our model achieves the highest accuracy, exhibiting the smallest performance drop of 1.4% OA from the ‘obj_nobg’ variant to the ‘obj_bg’ variant, surpassing all other networks in comparison.
TABLE XI: Robustness to background noise on ScanObjectNN. ‘obj_bg’, ‘obj_nobg’ stand for objects with and without noise.
@@
3DmFV[81] 69.8 63.0 6.8$\downarrow$
PointNet[24] 74.4 68.2 6.2$\downarrow$
PointNet++[25] 80.2 77.9 2.3$\downarrow$
SpiderCNN[12] 76.9 73.7 3.2$\downarrow$
DGCNN[13] 81.5 78.1 3.4$\downarrow$
PointCNN[36] 80.8 78.5 2.3$\downarrow$
SMTransformer(ours) 88.0 86.6 1.4normal-$\downarrow$

### v2

@@ IV-A Indoor Semantic Segmentation
TABLE IV: Semantic segmentation results on the SemanticKITTI. ‘*’ means the network does not use any extra data.
Year Methods
mIoU(%)
car
bicycle
motor.
truck
other-vehi.
person
bicyclist
motorcyc.
road
parking
sidewalk
other-gro.
building
fence
vegetation
trunk
terrain
pole
sign
2022 NIPS Point Transformer V2[36] 72.6
2023 ICCV RangFormer[15] 73.3 96.7 69.4 73.7 59.9 66.2 78.1 75.9 58.1 92.4 73.0 78.8 42.4 92.3 70.1 86.6 73.3 72.8 66.4 66.6
2024 CVPR Point Transformer V3[35]* 74.2
2024 TIM EPNet[5] 67.4
2025 IEEE ITJ BeyondPoints[22] 73.4
2025 TGRS CSFNet[23] 71.5 97.6 61.5 88.5 96.5 82.8 82.4 93.6 2.2 94.8 52.1 83.0 0 93.4 73.7 89.1 72.8 75.6 66.3 52.3
2025 TCSVT NUCNet[32] 73.6 97.4 70.9 74.9 60.2 63.6 79.2 77.6 54.6 91.6 71.5 77.7 45.5 92.8 70.9 86.9 75.3 72.6 66.9 68.9
Results: We compare our method with the recent state-of-the-art on S3DIS dataset, using three metrics i.e. mean class-wise intersection over union (mIoU), mean overall accuracy (mAcc), overall accuracy (OA) and parameters (Para.). To ensure a fair comparison with previous state-of-the-art point transformer methods, we reproduced their original code on our own workstation, maintaining the original conditions without using any extra data. Results are reported in Table I. Without the test time augmentation trick, our network demonstrates superior performance on three metrics i.e. 72.6% mIoU, 78.8% mAcc and 91.7% OA, respectively. It outperforms the DeepLA[41] by 0.7% mIoU in the term of mIoU. With the test time augmentation trick, our network still demonstrates superior performance on three metrics i.e. 73.6% mIoU, 79.0% mAcc and 91.8% OA. Compared to the Point Transformer V3, the performance of our method exceeds it by 0.5% mIoU, 0.7% mAcc and 0.7% OA, respectively, and our network parameters (i.e. 7.8 million) are reduced by about 68%. Fig. 5. shows visualizations of our results on S3DIS area 5 in comparison to Point Transformer V2. We can see that our method is more robust to the object boundaries.
The results on ScanNetV2 are illustrated in Table II. Similarly, to ensure a fair comparison, we reproduced the original code of several recent state-of-the-art point transformer methods on our own workstation without using any extra data. These methods include Stratified Transformer[16], Point Transformer V2 [35], ConDAForMer[6], Point Transformer V3[35] and Full Point Transformer [8]. Our network achieved the second-best result with a 76.2% mIoU, having the fewest parameters (7.8 million) in the compared transformer methods. Compared to the DeepLA, our method exhibits a substantial improvement of +2.1% mIoU. Although our method performs slightly worse than the Point Transformer V3, our network has only about 17% of its parameters.
IV-B Outdoor Semantic Segmentation
Datasets: We conduct experiments on two popular datasets: SWAN [12] and SemanticKITTI[2] dataset. The demanding SWAN dataset comprises 32 sequences of point clouds totalling 10,000 frames and containing approximately 0.9 billion points. Sequences 0 to 23 are allocated for training, while sequences 24 to 31 are designated for testing. The SemanticKITTI provides 22 sequence point clouds consisting of 43,552 frames. Adhering to standard practice, we employ sequences 0 to 10 (excluding 8) for training, use sequence 8 for validation and sequences 11 to 21 for testing. The labels for the test set are exclusively available to the online server, necessitating result submissions for remote evaluation.
Network Configurations: For semantic segmentation on SWAN, we opt not to employ voxelization to reduce point resolution; instead, we directly process the raw point cloud data. We set the maximum number of points to 80,000. We use the AdamW optimizer and set weight decay as 0.04. The base learning rate is set as 0.004 and the learning rate is scheduled by the MultiStepLR. We use the same data augmentation as the ones on SemanticKITTI to preprocess the input data. Our model undergoes training and testing phases with a batch size of 16 and 8 distributed across 4 GPUs. On SemanticKITTI, we set the voxel size as 5cm and the maximum number of voxels to 100,000. We use the AdamW optimizer and weight decay as 0.02. The base learning rate is set as 0.004 and the learning rate is scheduled by the Cosine. We adopt rotation, flip, scaling, and transformation to augment training data.
@@ IV-B Outdoor Semantic Segmentation
Results: We present the results for 20 classes of interest in the SWAN test frames in Table III. We compare our results to PointNet++ [36], PointConv [34], Cylinder3D [46], PolarNet [44], $\psi$-CNN [17], SAT3D [12], Point Transformer [45], Point Transformer V2 [36], and Point Transformer V3 [35]. These results are obtained by carefully training the models using the author-provided implements, following the original guidelines for hyper-parameter settings. As depicted in Table III, the mIoU values for these compared methods on the SWAN dataset are lower than those on the semanticKITTI dataset. This discrepancy can be attributed to the heightened complexity of the scenes in the SWAN dataset which was captured in dense central business district are of the city of Perth, Australia. Notably, on this dataset, our method demonstrates the best performance with a remarkable improvement of +4.2% in mIoU compared to the nearest competitor SAT3D. Besides, compared to the previous point transformers (e.g. Point Transformer V2 and V3), our method exceed them by 2.7% and 1.3% mIoU, respectively. Our method is generally able to show remarkable prediction accuracy for some small objects, including light poles, traffic sign and pedestrians.
Table IV presents the outcomes of our network alongside results from well-established methods on the SemanticKITTI dataset. Our approach demonstrates commendable performance, achieving 74.9% mIoU in these compared methods. Notably, compared to the cutting-edge projection-based method RangFormer, which is pre-trained on other datasets, our method exhibits superior performance (+1.6%) without any pre-training. Furthermore, our proposed method surpasses voxel partitioning and 3D convolution-based techniques, such as 2DPASS, and PVKD. Importantly, compared to the previous state-of-the-art methods, our method outperforms CSFNet and NUCNet by 3.4% mIoU and 1.3% mIoU, respectively. Our model showcases a remarkable understanding of certain small object categories, including poles, traffic signs and motorcyclists.
IV-C Ablation Studies
@@
We conduct ablation studies on the S3DIS dataset to demonstrate the effectiveness of SMTransformer, SAUB and Shared Point Position Encoding.
1) Effect of Various Components: Table V presents the quantitative results of our proposed modules (i.e., SMTB, SAUB, and SPE) in comparison with the Point Transformer (PT) series. Building on PT, Cases I–III progressively integrate SMTB, SAUB, and SPE. The results show that our design not only improves segmentation accuracy but also reduces computational overhead. In particular, Case III (the full model) achieves 73.6% mIoU with only 7.8M parameters, 2.65G FLOPs, and an inference time of 34ms, surpassing PT (70.6% mIoU, 7.8M parameters, 2.80G FLOPs, 36ms) in both accuracy and efficiency. Compared with PTV2 and PTV3, our model attains higher accuracy while requiring significantly fewer parameters (7.8M vs. 12.8M/46.19M) and shorter training time (16h vs. 22h/46.5h). The SPE makes the model lighter because transformer blocks within the same layer share a single position encoding, eliminating redundant parameters. Additionally, by providing a consistent positional representation, the model can focus on feature relationships rather than memorizing layer-specific positional biases. This design contributes to the model’s parameter efficiency and promotes a more unified geometric understanding.
TABLE V: Comparison of model efficiency based on the parameters (M), FLOPs (G), inference time (ms) and training time (h). Para.:Parameter. IT: Inference Time. $T$: Training Time.
Methods SMTB SAUB SPE mIoU Para. FLOPs IT $T$
PTV3 73.1 46.2 5.82 94 46.5
PTV2 71.6 12.8 3.64 42 22
PT 70.6 7.8 2.80 36 20
I 72.9 9.0 2.96 38 22
II 73.4 10.8 3.24 43 24
III 73.6 7.8 2.65 34 16
@@
2) Soft Mask: To demonstrate the effectiveness and versatility of the soft mask, we conduct a comparison with the hard mask. The hard mask, represented by binary mask, can be expressed as,
@@
3) Enhanced Position Encoding: We study the effects of different encoding strategies used for the position encoding in our Transformer encoder. We compare the proposed enhanced position encoding (EPE) with some classical position encoding strategies such as local position encoding (LPE) and global position encoding (GPE). For each of these cases, we test with learnable MLP or nonlearnable Sinusoidal[21] position encoding. The results are shown in Table VI. We can see that the performance of global position encoding is lower than that of local position encoding. The underlying cause is that global position encoding lacks the geometric connection information. When full position encoding strategy is incorporated into the Transformer, the network achieves the highest performance. Our results also show that the MLP encoder is more flexible than Sinusoidal encoder. This indicates our proposed point transformer with full position encoding has more geometric awareness.
TABLE VI: Segmentation performance of our model on S3DIS area 5 with different position encoding. The network does not include SADS block.
Encoder Strategy mIoU(%) Encoder Strategy mIoU(%)
Sinusoidal LPE 69.8 MLP LPE 70.2
Sinusoidal GPE 68.4 MLP GPE 69.3
Sinusoidal FPE 70.5 MLP FPE 71.5
4) Skip Attention-based Up-Sampling Block: To prove the effectiveness of our proposed skip attention-based up-sampling block (SAUB), we compare it with two types of up-sampling blocks including the transition up-sampling block (TUB)[45] and grid unpooling block (GUB)[36]. TUB consists of one projection layer, one interpolation, and one addition operation. GUB consists of one projection layer, one grid unpooling operation and one addition operation. The addition operation connects the features from the encoding and decoding layers. Table VII presents the performance of our network with various up-sampling blocks. The network employing the SAUB achieves the best performance with mIoU, mAcc, and OA values of 73.6%, 79.0%, and 91.8%, respectively, surpassing the performance of the network using the TUB by a significant margin.
TABLE VII: Segmentation performance of our model on S3DIS area 5 with different up-sampling blocks. TUB: transition up-sampling block, GUB: grid unpooling block, SAUB: skip-attention-based up-sampling block.
@@
SAUB(ours) 73.6 79.0 91.8
@@
IV-D Robustness Analysis
1) Robustness to Density: We compare the robustness of our model to inter- and intra-point cloud density with several typical baselines such as PointNet [25], PointNet++ [26], DGCNN[33], classical convolutional network such as PointConv [34], RS-CNN[20], DANet[9] and attention network PointASNL[40]. For a fair comparison, all the networks are trained on modelnet40_normal_resampled dataset[37] with 1024 points using only coordinates as the input. To showcase the robustness across inter-point clouds with varying densities, we utilize downsampled points of 512, 256, 128, and 64 as input to the trained model. To evaluate the robustness of intra-point cloud with various densities, we divide the 1024 points into four equal parts along the $X$ coordinate according to the point number, and then we randomly sample 128 points from each part in sequence. This generates the test samples with 896, 768, 640 and 512 points, respectively. The results are shown in Fig. 7. Our SMTransformer shows significantly superior robustness, surpassing existing approaches for both inter and intra-point cloud variations. In real-world automation applications, such as robotic perception in warehouses, point density varies drastically due to sensor range, occlusions, or motion. The superior density robustness of SMTransformer suggests that it can maintain reliable perception even under sparse or unevenly distributed point clouds, which is crucial for safety-critical automation tasks.
2) Robustness to Transformation: To demonstrate the robustness of our SMTransformer, we evaluate its performance on S3DIS and ModelNet40 under a variety of perturbations in the test data, including permutation, translation, scaling and jitter. As shown in Table VIII, on S3DIS, Point Transformer and Point Vector have a huge performance drop on scaling transformation. Our method exhibits remarkable stability across diverse transformations. Particularly noteworthy is its stable performance even amidst a 0.2 translation along the $X$, $Y$, and $Z$ axes and jitter. All methods are invariant to permutations. In terms of sensitivity to point scaling, SMTransformer performs relatively better when the scaling range is decreased. Our method achieves the best accuracy under all transformations on both segmentation and classification datasets. Robotic and automation systems frequently encounter geometric perturbations, such as sensor misalignment, scaling effects from varying distances, and translation shifts during motion. The stability of SMTransformer under these transformations indicates strong adaptability to such conditions, making it well-suited for robotic navigation and manipulation tasks where viewpoint and scale continuously change.
[…]
@@ IV-D Robustness Analysis
TABLE IX: Robustness to background noise on ScanObjectNN. ‘obj_bg’, ‘obj_nobg’ stand for objects with and without noise.
@@
[…]
3) Robustness to Noise: To assess the robustness of SMTransformer to noise, we conducted experiments using the PB_T50_RS variant of ScanObjectNN dataset, measuring the performance with and without background noise (denoted as ‘obj_bg’ and ‘obj_nobg’ respectively). Table IX presents a comparative analysis between our model and several baselines from [30]. We observe that the overall accuracy of all networks diminishes when trained and tested under conditions involving background noise. However, our model achieves the highest accuracy, exhibiting the smallest performance drop of 1.3% OA from the ‘obj_nobg’ variant to the ‘obj_bg’ variant, surpassing all other networks in comparison. In industrial and robotic environments, raw 3D scans often contain background clutter and measurement noise. The robustness of SMTransformer to noisy data demonstrates its potential for reliable operation in unstructured environments, such as factory floors or outdoor scenes, where clean point clouds cannot be guaranteed.